Operations Report – December 10th

Crew 200 Operations Report 10-Dec-2018

SOL: 8

Name of person filing report: Antoine Bocquier

Non-nominal systems: generator autostart
Notes on non-nominal systems: None

Generator (hours run): From 5:15 pm last night to 9:00 am. On tonight at 4:50 pm.

Solar— SOC 96%
Diesel Reading – 24 %
Propane Reading – 34 %
Ethanol Free Gasoline – N/A gallons

Water (auxillary tank) – 0 gallons
Water (static tank) – 380 gallons
Auxillary to Static tank transfer – no
Gallons transferred: 0
Water in GreenHab – refilled, high level
Water (loft) – Static to Loft Pump used – yes
Water Meter: 4

Toilet tank emptied: no

Deimos rover used: no, still not functional
Beginning charge:
Ending charge: 71%
Currently charging: no

Sojourner rover used: No
Beginning charge:
Ending charge: 100%
Currently charging: yes

Spirit rover used: no
Hours: 59.6
Beginning charge:
Ending charge: 100%
Currently charging: yes

Opportunity rover used: no
Hours: 41.1
Beginning charge:
Ending charge: 100%
Currently charging: yes

Curiosity rover used: no
Hours: 57.5
Beginning charge:
Ending charge: 100%
Currently charging: yes

Notes on rovers: Rovers were not used this week
# Hours the ATVs were Used today: 0
Notes on ATVs: ATVs were taken out of active service and are now parked away from the station

HabCar used and why, where? No
CrewCar used and why, where? No

General notes and comments: Eveything looks fine

Summary of internet: Nothing to report.

Summary of suits and radios: we performed an EVA this morning to study the radio coverage close to the Kissing Camel range. Spacesuit 3 does not seem to work, I am trying another charger

Summary of Hab operations: everything is fine, living, cooking and sleeping are still pleasant.

Summary of GreenHab operations: Regular activities

Summary of ScienceDome operations: We mopped the ground.

Summary of RAM operations: Nothing to report

Summary of any observatory issues: Musk Solar observatory – Nominal. Robotic Observatory – Was open all night and day long, we exchanged with Peter and CapCom to fix this and it finally closed around 5:00pm.

Summary of health and safety issues: security checks performed

Questions, concerns and requests to Mission Support: Nothing to report

Copyright © The Mars Society. All rights reserved. | Main Site